The weight 250 senior high school average 60 kg and the standard deviation is 6 kg how many children weigh between 50 kg and 75 kg


1
Expert's answer
2022-03-21T16:01:00-0400

"Z(50)=\frac{x-\mu}{\sigma}=\frac{50-60}{6}=-1.6667"

P(-1.6667<Z<0)=0.4525

"Z(75)=\frac{75-60}{6}=2.5"

P(0<Z<2.5)=0.4938

P(50<X<75)=0.4525+0.4938=0.9463

N(55<X<75)=0.9463x250=237